A catering service offers
6 appetizers,
4 main courses, and
8 desserts. A customer is to select
2 appetizers,
3 main courses, and
3 desserts for a banquet. In how many ways can this be done?

Found 2 solutions by Shin123, greenestamps:
Answer by Shin123(626)

There are 6C2+4C3+8C3 ways. There are 15+4+56=75 ways.

Answer by greenestamps(13367)



The answer is the PRODUCT of the numbers of ways to select the appetizers, main courses, and desserts -- not the SUM.

Correct answer: (6C2)*(4C3)*(8C3) = 15*4*56 = 3360